Examine the continuity of the function f:[1,4]→R defined by f(x)= [ x]+4/(3x-1)
1
Expert's answer
2020-03-09T11:27:30-0400

Note: we assume that "y=[x]" means "y=\lfloor x\rfloor". Which in turn means


"\lfloor x\rfloor=\left\{\left.m\in\mathbb{x}\right|m\le x\right\}"


( More information: https://en.wikipedia.org/wiki/Floor_and_ceiling_functions )

Since, "f : [1;4]\to\mathbb{R}" , then consider how the specified function "y=\lfloor x\rfloor" behaves on a given interval:


"\lfloor x\rfloor=\left\{\begin{array}{c}
1,\,\,\,\text{if}\,\,\,1\le x<2\\
2,\,\,\,\text{if}\,\,\,2\le x<3\\
3,\,\,\,\text{if}\,\,\,3\le x<4\\
4,\,\,\,\text{if}\,\,\,x=4
\end{array}\right."

Then, the function "f(x)" looks like this


"f(x)=\left\{\begin{array}{l}
\displaystyle\frac{5}{3x-1},\,\,\,\text{if}\,\,\,1\le x<2\\[0.5cm]
\displaystyle\frac{6}{3x-1},\,\,\,\text{if}\,\,\,2\le x<3\\[0.5cm]
\displaystyle\frac{7}{3x-1},\,\,\,\text{if}\,\,\,3\le x<4\\[0.5cm]
\displaystyle\frac{8}{3x-1},\,\,\,\text{if}\,\,\,x=4
\end{array}\right."

We see that it is necessary to study the continuity of the function at the points "x=2" , "x=3" ,

"x=4". For this, one-sided limits must be considered.

( More information: https://en.wikipedia.org/wiki/One-sided_limit )

In our case,


"\lim\limits_{x\to 1^+}f(x)=\lim\limits_{x\to 1^+}\frac{5}{3x-1}=\frac{5}{3\cdot 1-1}=\frac{5}{2}\\[0.5cm]
f(1)=\frac{5}{3\cdot 1-1}=\frac{5}{2}\\[0.5cm]
\boxed{\lim\limits_{x\to 1^+}f(x)=\frac{5}{2}=\frac{5}{2}=f(1)}\\[0.5cm]
\lim\limits_{x\to 2^-}f(x)=\lim\limits_{x\to 2^-}\frac{5}{3x-1}=\frac{5}{3\cdot 2-1}=1\\[0.5cm]
\lim\limits_{x\to 2^+}f(x)=\lim\limits_{x\to 2^+}\frac{6}{3x-1}=\frac{6}{3\cdot 2-1}=\frac{6}{5}\\[0.5cm]
\boxed{\lim\limits_{x\to 2^-}f(x)=1\neq\frac{6}{5}=\lim\limits_{x\to 2^+}f(x)}\\[0.5cm]
\lim\limits_{x\to 3^-}f(x)=\lim\limits_{x\to 3^-}\frac{6}{3x-1}=\frac{5}{3\cdot 3-1}=\frac{6}{8}\\[0.5cm]
\lim\limits_{x\to 3^+}f(x)=\lim\limits_{x\to 3^+}\frac{7}{3x-1}=\frac{7}{3\cdot 2-1}=\frac{7}{8}\\[0.5cm]
\boxed{\lim\limits_{x\to 3^-}f(x)=\frac{6}{8}\neq\frac{7}{8}=\lim\limits_{x\to 3^+}f(x)}\\[0.5cm]
\lim\limits_{x\to 4^-}f(x)=\lim\limits_{x\to 4^-}\frac{7}{3x-1}=\frac{7}{3\cdot 4-1}=\frac{7}{11}\\[0.5cm]
f(4)=\frac{8}{3\cdot4-1}=\frac{8}{11}\\[0.5cm]
\boxed{\lim\limits_{x\to 4^-}f(x)=\frac{7}{11}\neq\frac{8}{11}=f(x)}\\[0.5cm]"

Conclusion,

The function "f(x)" is discontinuous at the points "x=2" , "x=3" , "x=4" .

The function "f(x)" is continuous at the point "x=1" .